Chapter 21, Problem 024 Two tiny, spherical water drops, with identical charges of-5.71 × 10-16 C, have a center-to-center separation of 0.758 cm. (a) what is the magnitude of the electrostatic force acting between them? (b) How many excess electrons are on each drop, giving it its charge imbalance? (a) Number (b) Number Click if you would like to Show Work for this question: Open Show Work Units Units By accessing this Question Assistance, you will learn while you earn points based on the Point Potential Policy set by your instructor. Question Attempts: 0 of 5 used SAVE FOR LATER SUBMI
Answer #1

Solution :-

We know that

   Force   F   =   (1/4πε0)* q1 * q2 / r2

   a. F=9.0×10(-5.71× 10-16 )×(- 5.71×10-16 )/(0.758 × 10-2)2

         F   = 5.107×10-17  N( Repulsive)

   b.   By, quantization of charge principle

         q   =   n×e

         -5.71×10-16   =   n×(-1.6×10-19)

         n   = (5.71×10-16 / 1.6×10-19 ) = 3568.75

   Each drop has 3569 excess electrons